AnimeGhoul8863:

In your lab, a substance's temperature has been observed to follow the function T(x) = (x − 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Using complete sentences in your written answer, explain to your fellow scientists how to find the turning point of this function. Hint: The turning point of the graph is similar to the vertex of a quadratic function.

AnimeGhoul8863:

@Mehek

AnimeGhoul8863:

Im guessing T(x)=(x-4) * 3+6 since its next to the ()

AnimeGhoul8863:

so if im thinking correctly we first expand it? (x-4)*3+6=3(x-4)+6

AnimeGhoul8863:

3(x-4)=3x-12 which makes it 3x-12+6

AnimeGhoul8863:

Subtract the numbers -12 + 6 = -6 making it 3x-6

AnimeGhoul8863:

Tx=3x-6 we then subtract 3x from both sides Tx-3x=3x-6-3x

AnimeGhoul8863:

we then simplify Tx-3x=-6

AnimeGhoul8863:

then we factor Tx-3x=x(T-3)

AnimeGhoul8863:

x(T-3)=-6 Then divide by T-3;

AnimeGhoul8863:

\[\frac{ x(T-3) }{ T-3 }=\frac{ -6 }{ T-3 };\]

AnimeGhoul8863:

\[x=-\frac{ 6 }{ T-3}\]

AnimeGhoul8863:

\[T \neq 3\]

AnimeGhoul8863:

@Mehek am i correct?

Mehek:

Is it not T(x) = (x-4)^3 + 6?

AnimeGhoul8863:

Im stupid *smh* your right i misread it

AnimeGhoul8863:

UGHHHH i have to start over T~T

Mehek:

Which math are you taking

AnimeGhoul8863:

Algebra 2

AnimeGhoul8863:

(4,6) would be the vertex

Mehek:

Correct

AnimeGhoul8863:

So The Turning Point of this Equation would be (4,6) i say this because when graphing this Equation the Degree is 3 meaning its a cubic function meaning it increases but also decreases leaving the vertex to be (4,6)

AnimeGhoul8863:

@Mehek would this be correct?

Mehek:

Yes